# Capturing Stable HDR Videos Using a Dual-Camera System

### v1

@@ Abstract
In HDR video reconstruction, exposure fluctuations in reference images from alternating exposure methods often result in flickering. To address this issue, we propose a dual-camera system (DCS) for HDR video acquisition, where one camera is assigned to capture consistent reference sequences, while the other is assigned to capture non-reference sequences for information supplementation. To tackle the challenges posed by video data, we introduce an exposure-adaptive fusion network (EAFNet) to achieve more robust results. EAFNet introduced a pre-alignment subnetwork to explore the influence of exposure, selectively emphasizing the valuable features across different exposure levels. Then, the enhanced features are fused by the asymmetric cross-feature fusion subnetwork, which explores reference-dominated attention maps to improve image fusion by aligning cross-scale features and performing cross-feature fusion. Finally, the reconstruction subnetwork adopts a DWT-based multiscale architecture to reduce ghosting artifacts and refine features at different resolutions. Extensive experimental evaluations demonstrate that the proposed method achieves state-of-the-art performance on different datasets, validating the great potential of the DCS in HDR video reconstruction. The codes and data captured by DCS will be available at https://github.com/zqqqyu/DCS.
@@
HDR video, dual-camera system, multi-exposure image fusion
@@ I Introduction
Capturing the full range of illumination in high dynamic range (HDR) scenarios is a challenging task for standard digital cameras [1, 2]. Thanks to the multi-exposure fusion (MEF) technology [3, 4, 5], we can successfully capture HDR images in static scenarios by fusing multiple images of different exposure times. However, capturing HDR videos with a standard digital camera in dynamic scenes has not yet been effectively resolved. Early studies concentrated on designing hardware systems for directly capturing HDR video, including internal/external beam splitters [6, 7, 8], specialized sensors [9, 10], modulo cameras [11] and neuromorphic cameras [12], etc. Although these solutions show good potential for capturing high-quality HDR videos, their high cost and bulky design severely limit their widespread adoption.
Figure 1: Qualitative comparison between SAFNet and AE methods. Existing AE methods result in flickering artifacts (luminance fluctuations exceeding 30). In contrast, our system ensures consistent illumination across reference frames, enabling temporally consistent HDR reconstruction results (luminance fluctuations limited to under 1).
To address the limitations of these issues, researchers [13, 14, 15] have turned to computational methods that leverage inexpensive standard digital cameras, providing a promising direction for HDR video reconstruction. With the rapid advancement of deep learning, researchers make efforts to solve HDR imaging problems in dynamic scenes with deep neural networks. These methods can be broadly categorized into two types. One approach follows the MEF attempts, aiming to reconstruct an HDR image by fusing misaligned multi-exposure images. They typically use a specific exposure image as a reference frame for HDR reconstruction. As a result, capturing HDR video using these methods often requires a large number of additional exposure images, making it difficult to meet the real-time requirements of video capture. To overcome this limitation, alternating exposure (AE) methods [16, 17, 18, 19, 20] have been proposed, which allow varying exposures as reference frames and eliminate the need for additional exposure images. However, aligning the brightness of reference frames of different exposures and achieving consistent and stable lighting in the reconstruction video is still challenging. Although existing AE methods have achieved excellent performance on public benchmarks, they struggle to handle complex real-world scenes and lighting variations (shown in Fig. 1).
Generally, both existing MEF methods and AE methods struggle to achieve high-quality video capture in dynamic scenes. Compared to AE methods, MEF methods could produce HDR image sequences with stable and consistent lighting because of their use of a fixed-exposure reference. Noticing the wide use of multi-camera system in consumer electronic devices, it is possible to introduce multi-camera system for high-quality HDR video capture. Based on these insights, we develop a dual-camera system (DCS) for capturing high-quality HDR videos. In the DCS, one camera is assigned to capture reference images, while the other is assigned to capture non-reference images for MEF. In this way, the DCS could effectively mitigate the temporal inconsistencies associated with AE methods while avoiding the acquisition latency inherent in MEF.
Within DCS, HDR video reconstruction with MEF faces greater challenges than static images for two main reasons. First, exposures of input frames become uncertain. Second, the backgrounds from non-reference images are no longer aligned with the reference frame. Existing MEF methods can be categorized into two main approaches: optical-flow-based and attention-based. The former relies on optical flow to align inputs before fusion, which often suffers from a low accuracy in estimating optical flow. On the other hand, attention-based methods, while offering greater flexibility, mainly depend on the quality of the attention maps for fusion. Therefore, existing attention-based MEF methods encounter two main issues in reconstructing HDR videos within DCS. First, they fail to adequately leverage exposure information, relying primarily on texture and luminance to calculate attention maps, which may suffer from the various exposure settings. Second, they hardly consider the importance of reference frames, particularly in the computation of attention maps, where equal weights are often assigned to both reference and non-reference frames. This leads to attention being skewed by non-reference inputs, especially in complex, dynamic scenes, thereby hindering the effective elimination of artifacts.
Against the above limitations, we propose an exposure-adaptive fusion network (EAFNet) to achieve robust and generalized HDR reconstruction. Our EAFNet consists of a pre-alignment subnetwork, an asymmetric cross-feature fusion subnetwork, and a reconstruction subnetwork. The pre-alignment subnetwork consists of a global luminance alignment (GLA) and an exposure-guided feature selection module (EFSM) to leverage exposure information, enhancing meaningful details and ensuring that the most relevant features are preserved despite varying exposure levels. The asymmetric cross-feature fusion subnetwork aims to explore reference-dominated attention maps, align cross-scale features, and integrate cross-feature information to improve the fusion. Additionally, a multiscale reconstruction subnetwork is introduced to suppress ghosting artifacts and enhance fine details. Summary, our contributions can be summarized as follows:
@@
  • We propose a new HDR video generation paradigm that uses a fixed medium-exposure sequence as the temporal reference, while leveraging a second varying-exposure stream to enrich dynamic range. This design fundamentally improves temporal consistency and reconstruction stability.
  • We designed and implemented an asynchronous dual-camera system to validate the feasibility of our proposed paradigm and bridge the gap between algorithmic design and practical deployment. Furthermore, existing MEF methods can be seamlessly applied within our system, achieving temporally consistent reconstruction.
  • Built upon MEF methods, we propose a novel model design, EAFNet. The proposed pre-alignment subnetwork leverages exposure information and explores the intrinsic properties of the images, ensuring that the most relevant features receive attention despite varying exposure levels. An asymmetric cross-feature fusion subnetwork is introduced to explore reference-dominated attention maps, achieving smoother and more consistent fusion through cross-scale alignment and cross-feature interaction.
  • Our EAFNet achieves state-of-the-art results on public benchmarks. Both qualitative and quantitative experiments demonstrate the effectiveness of our EAFNet and DCS.
@@ II RELATED WORK
II-A1 Traditional Methods.
Traditional multi-exposure image fusion HDR imaging methods usually solve the ghosting problem in the image domain, and currently there are two main types of solutions:
Motion-aligned-based. These methods focus on a global alignment of the LDR input and then reject the unaligned pixels before image fusion to achieve de-ghosting. Bogoni et al. [4] used optical flow to estimate motion vectors between LDR images and warp the moving pixels accordingly. Kang et al. [16] improved the alignment results by using the exposure time information to transform the intensity of LDR images to the brightness domain before aligning them with the optical flow method. Sen et al. [5] proposed a patch-based energy minimization method that integrates alignment and HDR reconstruction into a joint optimization. Hu et al. [21] optimized alignment using luminance and gradient consistency in the transform domain. Hafer et al. [22] proposed an energy minimization approach to jointly compute HDR radiance and displacement fields. These motion-alignment methods heavily rely on the accuracy of the alignment algorithms.
Motion-rejection-based. Motion-rejection methods detect and exclude motion regions before fusion. Grosch et al. [23] identified motion regions based on intensity differences, while Jacobs et al. [13] used weighted variance to detect ghosting. He et al. [24] refined motion detection by combining joint probability density with graph-cut-based energy minimization. Zhang et al. [25] removed motion pixels using image gradient information, while Lee et al. [26] and Oh et al. [27] introduced rank minimization to detect motion and reconstruct HDR images. However, discarding motion regions often compromises information recovery in static areas.
II-A2 Deep Learning Approaches.
With the rapid advancement of deep learning, neural network-based methods for HDR image reconstruction have gained significant attention and have emerged as the prevailing approach in the field, these can be categorized primarily into the following types:
@@
Direct Feature Concatenation. Wu et al. [32] introduced a flow-free approach, treating HDR imaging as an image transformation problem. They first applied homography to align multiple frames globally and then used a multiscale network to map LDR images to HDR images directly. A similar approach was also employed in [33]. Niu et al. [34] and Li et al. [35] used Generative Adversarial Networks (GANs) for HDR reconstruction from multi-exposure LDR images with significant motion. However, they still face challenges in handling extreme motion and dynamic scenes.
Attention-based Alignment. AHDRNet [36] introduced the attention mechanism into multi-exposure HDR reconstruction for dynamic scenes, coupling it with a densely connected dilated residual module to achieve alignment and fusion. Liu et al. [37] introduced a Pyramid Cascading Deformable (PCD) module, using multiscale deformable convolutions for feature alignment in LDR images. Chen et al. [38] proposed novel attention mechanisms and progressive texture fusion to improve reconstruction. DomainPlus [39] utilized multiscale attention with fronted fusion to perform coarse-to-fine fusion. HDR-Trans et al. [40] added Transformer to the CNN-based HDR architecture, combining the advantages of CNN in extracting local information and Transformer in capturing global information. Chen et al. [41] proposed a transformer model for HDR imaging with three steps: alignment, fusion, and reconstruction. Yan et al. [42] proposed a HyHDRNet consisting of a content alignment subnetwork and a transformer-based fusion subnetwork to improve performance. SCTNet [43] integrated spatial and channel attention in a transformer-based network to improve semantic coherence. Despite advances in these approaches, they still generate undesirable ghosts in high-contrast regions.
II-A3 Dual-lens-based system
Dual-lens HDR imaging leverages paired images with different exposures to synthesize high-quality HDR content. Dong et al. [44] employ a soft-warping step to align long-exposure features to the short-exposure reference, Li et al. [45] propose a dual-stage model: a guided 3D exposure correction network estimates a coarse HDR image from the short exposure input, followed by a guided denoising transformer that refines the result using long-exposure features. While effective under controlled conditions, both frameworks assume the only misalignment present is 1D parallax between image pairs. This assumption is embedded in their data collection and rectification procedures, simplifying alignment to left–right disparities but limiting their ability to model more complex distortions. In dynamic scenes, exposure time mismatch introduces additional misalignments: high-exposure frames exhibit motion blur not present in low-exposure frames.
@@ II-B HDR Video Reconstruction
Traditional Methods. Similar to multi-exposure HDR image fusion, most methods utilize alternating exposure sequences to align multiple frames for HDR video reconstruction. Kang et al. [16] introduced the first algorithm in this category, employing global and local registration to align multi-exposure images at each time step and merge them into a single HDR frame. Kalantari et al. [17] employed patch-based optimization to reconstruct missing images from different exposures. Mangiat et al. [17] applied block-based motion estimation with refined motion vectors and later improved this method with HDR filtering to eliminate block artifacts [47]. Additionally, Gryaditskaya et al. [18] proposed an adaptive weighting algorithm to minimize motion artifacts. Li et al. [48] proposed a statistical method to avoid exact alignment estimation. However, these methods are typically time-consuming and often introduce visible artifacts.
Deep Learning Methods. Kalantari et al. [49] proposed the first deep learning-based method for generating HDR video from alternating exposure sequences. Chen et al. [19] developed a two-stage coarse-to-fine framework that integrated deformable convolutions with HDR video reconstruction. LAN-HDR [50] aligned adjacent frames with the reference frame using luminance-based attention scores. HDRFlow [51] designed an efficient flow network with multiscale large kernels to handle large motions. Cui et al. [20] introduced exposure completion for neural HDR rendering, using feature interpolation and a novel decoding architecture to improve HDR video reconstruction. Nonetheless, the varying exposures of the reference frames create a significant challenge in maintaining temporal consistency in video reconstruction.
@@
III DUAL-CAMERA SYSTEM FOR HDR VIDEO
Overview. A mainstream solution for HDR video reconstruction is generating video from image sequences captured at alternating exposure levels. However, this approach inherits the exposure limitations of the reference frame, often resulting in visible artifacts and temporal flickering in the generated video, significantly constraining its practical application. To address this, we propose a dual-camera system (As shown in Fig. 2 (b)), where the primary camera continuously captures medium-exposure images to ensure temporal consistency, while the secondary camera captures low-exposure and high-exposure images to provide complementary information.

### v2

@@ Abstract
High Dynamic Range (HDR) video acquisition using the alternating exposure (AE) paradigm has garnered significant attention due to its cost-effectiveness with a single consumer camera. However, despite progress driven by deep neural networks, these methods remain prone to temporal flicker in real-world applications due to inter-frame exposure inconsistencies. To address this challenge while maintaining the cost-effectiveness of the AE paradigm, we propose a novel learning-based HDR video generation solution. Specifically, we propose a dual-stream HDR video generation paradigm that decouples temporal luminance anchoring from exposure-variant detail reconstruction, overcoming the inherent limitations of the AE paradigm. To support this, we design an asynchronous dual-camera system (DCS), which enables independent exposure control across two cameras, eliminating the need for synchronization typically required in traditional multi-camera setups. Furthermore, an exposure-adaptive fusion network (EAFNet) is formulated for the DCS system. EAFNet integrates a pre-alignment subnetwork that aligns features across varying exposures, ensuring robust feature extraction for subsequent fusion, an asymmetric cross-feature fusion subnetwork that emphasizes reference-based attention to effectively merge these features across exposures, and a reconstruction subnetwork to mitigate ghosting artifacts and preserve fine details. Extensive experimental evaluations demonstrate that the proposed method achieves state-of-the-art performance across various datasets, showing the remarkable potential of our solution in HDR video reconstruction. The codes and data captured by DCS will be available at https://zqqqyu.github.io/DCS-HDR/.
@@
HDR video, multi-exposure image fusion, luminance alignment
@@ I Introduction
Capturing the full range of illumination in high dynamic range (HDR) scenarios is a challenging task for standard digital cameras [10, 18, 29]. Thanks to the multi-exposure fusion (MEF) technology [1, 45], we can successfully capture HDR images in static scenarios by fusing multiple images of different exposure times. However, capturing HDR videos with a standard digital camera in dynamic scenes has not yet been effectively resolved [51, 50]. Early studies concentrated on designing hardware systems for directly capturing HDR video, including internal/external beam splitters [28, 39, 47], specialized sensors [7, 15, 20, 38], modular cameras [61] and neuromorphic cameras [14], etc. Although these solutions show good potential for capturing high-quality HDR videos, their high cost and bulky design severely limit their widespread adoption.
Figure 1: Overview of challenges in the alternating exposure (AE) paradigm and our proposed dual-stream paradigm for stable HDR video generation. (a) existing AE paradigm fuses alternating exposures (LE/ME/HE) with varying reference frames, resulting in flickering artifacts (luminance fluctuations exceeding 30). (b-d) Our dual-stream paradigm employs a fixed-exposure reference sequence for stable luminance anchoring, while a separate stream handles exposure-variant enhancement, ensuring consistent illumination across reference frames, enabling temporally consistent HDR reconstruction results (luminance fluctuations limited to under 1).
The alternating exposure (AE) paradigm, first introduced by Kang et al. [25], enables low-cost HDR video acquisition using a single camera with rapidly switching exposures, but it also introduces challenges in maintaining temporal consistency and suppressing ghosting artifacts. Traditional AE-based methods adopt strategies such as motion-compensated fusion [24, 13] and optical flow-based reconstruction [35] to align and merge complementary exposures for HDR reconstruction, but they remain highly sensitive to motion misalignment, struggle with occlusions and non-rigid deformations, and often rely on hand-crafted heuristics that lack generalization. Recent deep learning approaches [21, 9, 54] have advanced this direction by employing deep neural architectures with stronger feature representations and learning-based alignment mechanisms. These methods have achieved promising results by training on synthetic datasets with well-structured exposure patterns and often produce temporally consistent outputs under controlled settings. However, in real-world scenarios, variations in brightness distribution caused by changes in scene content, illumination, or sensor response can disrupt the learned alignment and fusion behavior. As a result, networks that perform well in ideal conditions often suffer from temporal flickering or inconsistent reconstruction when deployed in real environments (see Fig. 1(d)).
This instability highlights a fundamental limitation of the AE paradigm itself, rather than just model design, thus leading us to explore a new solution that circumvents inter-frame exposure inconsistency. Specifically, in the AE paradigm, detail selection and global luminance determination are inherently entangled at the input level. Existing methods [9, 54] typically process both aspects within a single modeling stream, without explicit disentanglement in input design or feature representation. As a result, the network must simultaneously decide which regions to source from each exposure and how to establish the overall luminance mapping within the same representational space. As shown in Fig. 1, this tight coupling leads to feature interference and makes the learned mapping strongly dependent on the exposure patterns seen during training, thereby reducing robustness and temporal stability under distribution shifts. To address this, we propose a dual-stream generation paradigm that decouples temporal luminance anchoring from exposure-variant detail reconstruction, thereby reducing feature interference and enhancing generalization across varying exposure patterns and scene conditions. Our paradigm is meticulously designed at both the hardware system and algorithmic levels, forming a complete HDR video generation solution that supports practical deployment and robust reconstruction in real-world environments.
Our hardware architecture is designed to implement an asynchronous dual-camera system (DCS) that capitalizes on the prevalent availability of multi-camera hardware in consumer electronic devices. Unlike prior dual-camera HDR imaging systems [11, 31] designed for static scenes under strict synchronization constraints, our DCS is tailored for dynamic video scenarios. By relaxing the synchronization requirement, it supports independent exposure control across two cameras. This design not only facilitates flexible and practical deployment without precision alignment but also avoids the frame rate bottleneck imposed by long-exposure frames in synchronized systems. As a result, our system enables high-quality HDR video capture at high frame rates. Specifically, the designed DCS system utilizes a fixed-exposure reference sequence as the temporal baseline, while simultaneously leveraging a varying-exposure stream to enrich the dynamic range. This system-level approach addresses key limitations of existing video methods. The reference stream mitigates the temporal inconsistencies and flickering commonly observed in AE-based single-camera systems, while the division of labor between the two cameras eliminates the need for precise synchronization inherent in traditional multi-camera systems, allowing for independent and flexible exposure control. Furthermore, the proposed DCS remains fully compatible with most existing image HDR deghosting algorithms, enabling temporally consistent and robust HDR video reconstruction.
Building on the proposed paradigm and DCS system, we now turn to the algorithmic level. We first revisit existing HDR deghosting approaches to extract insights and expose their limitations. Current methods can be broadly divided into optical-flow-based [44, 42] and attention-based [56, 34] strategies. The former relies on optical flow to align inputs before fusion, which often suffers from low accuracy in estimating optical flow. On the other hand, attention-based methods, while offering greater flexibility, their attention maps are typically computed only from texture and luminance cues, without explicitly leveraging exposure information. Moreover, they generally assign equal importance to reference and non-reference frames, which allows attention to be skewed by non-reference inputs in complex scenes. In our DCS, the presence of parallax introduces additional variations between inputs, making it even more important to design attention mechanisms that emphasize reference frames. Consequently, both categories exhibit limited robustness in dynamic, exposure-varying, and parallax-affected settings, where inaccurate alignment and exposure-insensitive attention lead to ghosting and color distortion. To address this, we propose an exposure-adaptive fusion network (EAFNet) tailored for the proposed DCS system to achieve robust and generalized HDR reconstruction. Our EAFNet consists of a pre-alignment subnetwork, an asymmetric cross-feature fusion subnetwork, and a reconstruction subnetwork. The pre-alignment subnetwork consists of a global luminance alignment (GLA) and an exposure-guided feature selection module (EFSM) to leverage exposure information, enhancing meaningful details and ensuring that the most relevant features are preserved despite varying exposure levels. The asymmetric cross-feature fusion subnetwork aims to explore reference-dominated attention maps, align cross-scale features, and integrate cross-feature information to improve the fusion. Additionally, a reconstruction subnetwork is introduced to suppress ghosting artifacts and enhance fine details.
In summary, our contributions can be summarized as follows:
@@
  • Paradigm: We introduce a dual-stream HDR video generation paradigm that explicitly decouples temporal luminance anchoring from exposure-variant detail reconstruction. Our approach employs a fixed-exposure stream to maintain temporal alignment across frames, while a complementary stream with varying exposures enhances the dynamic range. This design fundamentally improves temporal consistency and reconstruction stability.
  • System: We design and implement an asynchronous dual-camera system to validate the feasibility of our proposed solution and bridge the gap between algorithmic design and practical deployment. Unlike traditional synchronized setups constrained by long-exposure frames, our system enables high-frame-rate video capture in dynamic scenes by supporting independent exposure control without requiring hardware-level synchronization. Moreover, the system seamlessly integrates with existing image deghosting methods to achieve temporally consistent video reconstruction.
  • Method: To support our dual-camera system, we propose a novel model design, EAFNet. We leverage exposure information and explore the intrinsic properties of the images, ensuring that the most relevant features receive attention under different exposure conditions. Furthermore, we explore reference-dominated attention and effectively fuse complementary features across exposures, enabling robust representation under diverse lighting scenarios.
  • Results: Our proposed dual-stream HDR video generation paradigm demonstrates significant advancements across multiple aspects. Through extensive qualitative and quantitative experiments, we validate the effectiveness of our contribution.
The rest of this article is organized as follows. Section II presents the related works. Section III introduces our solution for video HDR capturing and our dual-camera system. Section IV introduces the proposed EAFNet in detail. Section V presents the experimental results, and the conclusion is given in Section VI.
@@ II RELATED WORK
II-A1 Traditional Methods
Traditional multi-exposure HDR image fusion methods address the ghosting problem using two main approaches: motion-aligned-based and motion-rejection-based methods. Motion-aligned methods focus on globally aligning LDR images, followed by rejection of unaligned pixels to eliminate ghosting. Early works by Bogoni et al. [1] used optical flow for motion estimation, while later improvements by Kang et al. [25] incorporated exposure time information for enhanced alignment. Further advancements include patch-based energy minimization by Sen et al. [45] and luminance-based optimization by Hu et al. [17]. Motion-rejection methods, on the other hand, identify and exclude motion regions before fusion. Grosch et al. [12] and Jacobs et al. [19] detected motion using intensity differences and weighted variance, respectively, while He et al. [16] and Zhang et al. [59] refined motion detection with graph-cut and image gradient techniques.
II-A2 Deep Learning Approaches
With the rapid advancement of deep learning, neural network-based methods for HDR image reconstruction have gained significant attention and have emerged as the prevailing approach in the field. These can be categorized primarily into the following types:
@@ II-A2 Deep Learning Approaches
Direct Feature Concatenation. Wu et al. [53] introduced a flow-free approach, treating HDR imaging as an image transformation problem. A similar approach was also employed in [57]. Niu et al. [41] and Li et al. [30] used Generative Adversarial Networks (GANs) for HDR reconstruction from multi-exposure LDR images with significant motion. However, they still face challenges in handling extreme motion and dynamic scenes.
Attention-based Alignment. AHDRNet [56] first introduced the attention mechanism into multi-exposure HDR reconstruction for dynamic scenes, coupling it with a densely connected dilated residual module to achieve alignment and fusion. Subsequent works focused on refining alignment strategies. Liu et al. [33] introduced a Pyramid Cascading Deformable (PCD) module, using multi-scale deformable convolutions for feature alignment in LDR images. Chen et al. [3] and DomainPlus [62] further improved fusion through progressive attention and coarse-to-fine multi-scale designs. Transformer-based methods such as HDR-Trans [34], Chen et al. [4], and HyHDRNet [55] integrated global context modeling into HDR pipelines, enhancing feature representation. SCTNet [46] combined spatial and channel attention to improve semantic consistency. Despite advances in these approaches, most existing methods lack explicit modeling of exposure variations, resulting in sensitivity to brightness variations. Moreover, treating reference and non-reference frames equally during fusion often leads to ghosting artifacts, particularly in high-contrast regions dominated by non-reference content.
Dual-lens-based system. Dual-lens HDR imaging leverages paired images with different exposures to synthesize high-quality HDR content. Dong et al. [11] and Li et al. [31] align features under the assumption of simple 1D parallax between rectified stereo pairs. While effective under controlled conditions, both frameworks assume the only misalignment present is 1D parallax between image pairs. This assumption is embedded in their data collection and rectification procedures, simplifying alignment to left–right disparities but limiting their ability to model more complex distortions. In dynamic scenes, mismatched exposure times introduce complex misalignments, including motion blur in long-exposure frames that cannot be captured by disparity-based alignment alone.
@@ II-B HDR Video Reconstruction
Traditional Methods. Most HDR video reconstruction methods utilize alternating exposure sequences to align multiple frames for HDR video reconstruction. Kang et al. [25] introduced the first algorithm in this category, employing global and local registration to align multi-exposure images at each time step and merge them into a single HDR frame. Mangiat et al. [35] applied block-based motion estimation with refined motion vectors and later improved this method with HDR filtering to eliminate block artifacts [36]. Additionally, Gryaditskaya et al. [13] proposed an adaptive weighting algorithm to minimize motion artifacts. Li et al. [32] proposed a statistical method to avoid exact alignment estimation. However, these methods are typically time-consuming and often introduce visible artifacts.
Deep Learning Methods. Kalantari et al. [21] proposed the first deep learning-based method for generating HDR video from alternating exposure sequences. Chen et al. [2] developed a two-stage coarse-to-fine framework that integrated deformable convolutions with HDR video reconstruction. LAN-HDR [8] aligned adjacent frames with the reference frame using luminance-based attention scores. HDRFlow [54] designed an efficient flow network with multi-scale large kernels to handle large motions. Cui et al. [9] introduced exposure completion for neural HDR rendering, using feature interpolation and a novel decoding architecture to improve HDR video reconstruction. Nonetheless, alternating exposure setups inherently introduce brightness inconsistencies across reference frames, leading to temporal instability in the reconstructed video. In contrast, our design adopts a fixed-exposure reference strategy, which naturally preserves brightness consistency and mitigates flickering artifacts.
@@
III DUAL-CAMERA SYSTEM FOR CAPTURING HDR VIDEO
Overview. An inexpensive mainstream solution for HDR video reconstruction is generating video from image sequences captured at alternating exposure levels [21, 9, 54]. However, this approach inherits the exposure limitations of the reference frame, often resulting in visible artifacts and temporal flickering in the generated video, significantly constraining its practical application. To address this, we propose a dual-stream paradigm supported by a dual-camera system (as shown in Fig. 2), where the primary camera continuously captures medium-exposure frames to ensure temporal consistency, while the secondary camera captures low-exposure and high-exposure frames to provide complementary information.